OpenStudy (anonymous):

anti-derivative of \[2\pi \cos \pi t\]

OpenStudy (hba):

Just take the Integral.

OpenStudy (anonymous):

is it \[2\pi \sin \pi t\]

OpenStudy (anonymous):

o wait nvm forgot about the integral rules for trig functions